# On finding optimal collective variables for complex systems by minimizing the deviation between effective and full dynamics

### v1

@@ 2.2 Spectrum and timescales
In this section we study the spectrum of $\mathcal{T}$ as well as its implications on the speed of convergence of the dynamics $X_{n}$ to equilibrium.
We assume that $X_{n}$ is reversible and $\mathcal{T}$ is self-adjoint with respect to the inner product (3) (see (9)), such that its spectrum $\sigma(\mathcal{T})$ is real. To simplify the discussion, we further confine ourself to the case where each nonzero element of $\sigma(\mathcal{T})$ is an eigenvalue $\lambda\in\mathbb{R}$, for which the (integral) equation
@@
Proof.
To prove the first claim, let us apply [40, Theorem VI.23] and show that $\mathcal{T}$ is a Hilbert-Schmidt operator if (20) holds. Obviously, by the definition (2) we have
Therefore, the conditions of [40, Theorem VI.23] hold and it follows that $\mathcal{T}$ is a Hilbert-Schmidt operator. This in turn implies that $\mathcal{T}$ is compact and any nonzero element in its spectrum $\sigma(\mathcal{T})$ is an eigenvalue (see [40, Theorem VI.22] and [40, Theorem VI.15]).
Next, we prove the second claim. Assume without loss of generosity that the eigenfunction $\varphi$ corresponding to the eigenvalue $\lambda$ is normalized such that $\langle\varphi,\varphi\rangle_{\mu}=1$. Multiplying both sides of (19) by $\varphi$ and integrating with respect to $\mu$, we get $\lambda=\langle\mathcal{T}\varphi,\varphi\rangle_{\mu}$. Concerning the upper bound in the second claim, using (11)–(12) we can derive that $0\leq\mathcal{E}(\varphi)=\langle(I-\mathcal{T})\varphi,\varphi\rangle_{\mu}=1-\lambda$, which implies $\lambda\leq 1$. If $\lambda=1$, we must have $\mathcal{E}(\varphi)=0$ and therefore (11) implies that $\varphi$ is constant, since both $p$ and $\pi$ are positive. For the lower bound, similar to the proof of (12) in Lemma 13, we have (since $\mathcal{T}$ is self-adjoint)
$0\leq\frac{1}{2}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}p(x,y)(\varphi(y)+\varphi(x))^{2}\pi(x)\,\mathrm{d}{x}\mathrm{d}{y}=\langle(I+\mathcal{T})\varphi,\varphi\rangle_{\mu}=1+\lambda\,,$
The next proposition gives conditions which guarantee that all eigenvalues are nonnegative.
@@
    There is a unique self-adjoint linear operator $\mathcal{A}$ on $\mathcal{H}$ such that $\mathcal{T}=\mathcal{A}^{2}$.
@@
Note that by definition
$\langle\mathcal{T}f,f\rangle_{\mu}=\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}p(x,y)f(y)f(x)\pi(x)\,\mathrm{d}{x}=\mathbb{E}_{X_{0}\sim\mu}(f(X_{0})f(X_{1}))\,,\quad f\in\mathcal{H}\,.$ (22)
Therefore, the equivalence between (a) and (b) follows from the variational principle for the smallest eigenvalue of $\mathcal{T}$. To prove the equivalence between (b) and (c), we notice that (22) implies that (b) is equivalent to $\mathcal{T}\geq 0$. Therefore, applying the square root lemma [40, Theorem VI.9] we see that (b) implies (c). Conversely, assuming (c), we have $\langle\mathcal{T}f,f\rangle_{\mu}=\langle\mathcal{A}^{2}f,f\rangle_{\mu}=\langle\mathcal{A}f,\mathcal{A}f\rangle_{\mu}\geq 0$ and therefore (b) holds. This shows that (b) and (c) are equivalent. ∎
@@
Roughly speaking, the condition (b) of Proposition 2 describes that “in the average sense, $f(X_{n})$ does not change sign in two consecutive steps”. Intuitively, the equivalence between (a) and (b) of Proposition 2 tells that the eigenvalues of $\mathcal{T}$ are all nonnegative if and only if the underlying Markov chain makes minor moves in a single step.
A situation under which the condition (c) is met is when the chain $X_{n}$ can be embedded into another Markov process $Y_{m}$, such that $X_{n}=Y_{2n}$ for all $n\geq 0$. In particular, suppose that $X_{n}$ describes the states of a continuous-in-time reversible diffusion process governed by a SDE with a generator $\mathcal{L}$ and let $\tau>0$ be the lag-time (see the first case in Example 2.1). Then, $\mathcal{T}=\mathrm{e}^{\tau\mathcal{L}}=\mathcal{A}^{2}$ with $\mathcal{A}=\mathrm{e}^{\frac{\tau}{2}\mathcal{L}}$. In fact, in this case all eigenvalues of $\mathcal{T}$ are positive.
In general, it is possible that a transfer operator $\mathcal{T}$ has negative eigenvalues $\lambda\in(-1,0)$ under the assumptions in Proposition 1. Such negative eigenvalues (and the corresponding eigenfunctions) may even be important in studying the large-time dynamics of $X_{n}$, if they are close to $-1$. Similar to Theorem 2.2 below, variational characterization for these negative eigenvalues (which are closest to $-1$) can be obtained and applied to developing numerical methods. However, we will not consider negative eigenvalues, since we are mainly interested in the case where the Markov chains come from reversible SDEs and the eigenvalues in this case are all positive according to the discussion above.
@@
Moreover, by Hilbert-Schmidt theorem [40, Theorem VI.16], we can choose the corresponding eigenfunctions $(\varphi_{i})_{i\geq 0}$, with $\varphi_{0}\equiv 1$, such that they form an orthonormal basis of $\mathcal{H}$. For a function $f\in\mathcal{H}$, from (2) we obtain
@@
where $\mathbb{E}_{\mu}(f)$ denotes the mean value of $f$ under $\mu$. Therefore, as the Markov chain evolves, the expectation $\mathbb{E}(f(X_{n})|X_{0}=x)$ converges to $\mathbb{E}_{\mu}(f)$ (i.e. the value at equilibrium), and the speed of convergence is determined by the leading eigenvalues of $\mathcal{T}$ that are closest to $1$. Estimating these leading eigenvalues and computing the corresponding eigenfunctions help to understand the long-time dynamical behaviors (e.g. identifying large timescales and metastable states) of the process $X_{n}$. In this regard, let us quote from [57] the following variational characterization of the large eigenvalues and the corresponding eigenfunctions of $\mathcal{T}$. We also refer to [56, 55] for similar variational characterizations using generators, whose proofs can be adapted to prove Theorem 2.2 after minor modifications.
@@
$\sum_{i=1}^{m}\omega_{i}(1-\lambda_{i})=\min_{f_{1},\dots,f_{m}}\sum_{i=1}^{m}\omega_{i}\mathcal{E}(f_{i})\,,$ (25)
@@
$\mathbb{E}_{\mu}(f_{i})=0,$ $\quad\forall\,i\,\in\{1,\dots,m\}\,,$ (26)
$\langle f_{i},f_{j}\rangle_{\mu}=\delta_{ij}\,,$ $\quad\forall\,i,j\,\in\{1,\dots,m\}\,.$
@@
In the following remark we discuss the connection between Theorem 2.2 and the variational formulations in VAMP [50]. Despite of the similarities, we emphasize that the variational formulation in (25) (and also in (28) in the remark below) has the advantage in that it results in a minimization task rather than a maximization task, which is more convenient for numerical computations (see [57, 27] for loss functions in deep learning based on Theorem 2.2, and see [6] for an alternative approach). More importantly, by employing the Dirichlet energy $\mathcal{E}$, we are able to draw a connection between eigenfunctions and slow CVs using Theorem 2.2 (see Theorem 2.3 and the discussions after Remark 2.3), and obtain a unified framework that characterizes both the spectrum and transition rates (see Proposition 4 in Section 2.3) of the process.
@@
Theorem 2.2 is similar to the variational formulation in VAMP in the reversible case (see [50, Appendix D] and [52]). In fact, using the identity (12) in Lemma 13 and the normalization constraints in (26), we can write (25) as
$\sum_{i=1}^{m}\omega_{i}\lambda_{i}=\max_{f_{1},\dots,f_{m}}\sum_{i=1}^{m}\omega_{i}\langle f_{i},\mathcal{T}f_{i}\rangle_{\mu}\,,$ (27)
which recovers the VAMP-1 score defined in [50] when $\omega_{i}=1$ for $1\leq i\leq m$ (the eigenvalue $\lambda_{0}=1$ and its eigenfunction are excluded by the first constraint in (26)). Similarly, to make a connection to the VAMP-r score in [50] for an integer $r>1$, we note that the same proof of Theorem 2.2 implies that
$\sum_{i=1}^{m}\omega_{i}(1-\lambda_{i}^{r})=\min_{f_{1},\dots,f_{m}}\sum_{i=1}^{m}\omega_{i}\mathcal{E}^{(r)}(f_{i})$ (28)
holds under the constraints (26), where
$\mathcal{E}^{(r)}(f)=\frac{1}{2}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}p^{(r)}(x,y)(f(y)-f(x))^{2}\pi(x)\,\mathrm{d}{x}\,\mathrm{d}{y}$
denotes the Dirichlet energy corresponding to the transition density $p^{(r)}(x,y)$ that describes jumps from $x=X_{n}$ to $y=X_{n+r}$ (note that $(\lambda_{i}^{r})_{i\geq 0}$ are the eigenvalues of the transfer operator obtained with the lag-time $r\tau$). The formulation (28) recovers the VAMP-r score in [50] by the same reasoning used to derive (27). The Dirichlet energies in both (25) and (28) can be estimated using trajectory data of the underlying dynamics (see (13) in Lemma 13 and discussions in [57]).
Finally, we note that choices on different VAMP-r scores and different sizes of lag-times are discussed in [31, 6]. In fact, the identity (28) and the fact that $p^{(r)}$ is the transition density for jumps in $r$ steps imply that choosing VAMP-r scores for different integers $r$ with fixed lag-time $\tau$ is (at the theoretical level) equivalent to choosing lag-time $r\tau$ for different $r$ with fixed VAMP-1 score. In this sense, it suffices to tune one of these two quantities in practice while keeping the other one fixed.
@@
For simplicity we first assume $k=1$ and consider a scalar CV map $\xi:\mathbb{R}^{d}\rightarrow\mathbb{R}$. We rely on the observation that a good CV map $\xi$ that is capable of indicating the occurrences of essential transitions of $X_{n}$ on large timescales is necessarily capable of filtering out the noisy fluctuations on the smallest timescales. In other words, $\xi$ should be a slow variable such that the observations $\xi(X_{0}),\xi(X_{1}),\dots$ vary slowly (in most of the time unless essential transitions occur) as the dynamics $X_{n}$ evolves. This reasoning suggests that a natural way to find a good CV map $\xi$ is by minimizing the average of squared variations
@@
The above reasoning can be extended to the case where $m\geq 1$. Let $|z|_{\omega}^{2}:=(\sum_{i=1}^{m}\omega_{i}|z_{i}|^{2})^{1/2}$ denote the weighted $l^{2}$ norm of vectors $z=(z_{1},\dots,z_{m})^{\top}\in\mathbb{R}^{m}$, where the weights $\omega=(\omega_{1},\dots,\omega_{m})^{\top}$ are assumed to be both positive and non-increasing, i.e. $\omega_{1}\geq\omega_{2}\geq\dots\geq\omega_{m}>0$. The following result characterizes the leading eigenfunctions as slow variables.
@@
The CV map $\xi=(\varphi_{1},\dots,\varphi_{m})^{\top}$ defined by the first eigenfunctions $\varphi_{1},\dots,\varphi_{m}$ minimizes the average of squared variations
@@
among functions $f=(f_{1},\dots,f_{m})^{\top}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ whose components satisfy (26).
@@
The conclusion follows directly from Lemma 13 and Theorem 2.2. ∎
Note that the norm in (30) is an extension of the norm in (29) when $k>1$. The weights $\omega$ are introduced in order to remove the non-uniqueness of the minimizer of (30) due to reorderings of $f_{1},f_{2},\dots,f_{m}$. Condition (26) guarantees that $f_{1},f_{2},\dots,f_{m}$ are non-constant, normalized, and pairwise orthogonal (therefore linearly independent).
@@ 2.3 Transition rates
In this section, we study the transition of the process $X_{n}$ between two subsets $A,B\subset\mathbb{R}^{d}$ in the framework of TPT [46, 9, 10]. Although TPT has been extended to different types of Markov chains [32, 48, 15, 14], here we present new results on the transition rates of Markov processes that are discrete-in-time and continuous-in-space.
Let us assume that $A,B\subset\mathbb{R}^{d}$ are two closed subsets with smooth boundaries $\partial A$ and $\partial B$, respectively, such that $A\cap B=\emptyset$ and the complement of their union is nonempty, i.e. $(A\cup B)^{c}:=\mathbb{R}^{d}\setminus(A\cup B)\neq\emptyset$. Recall that TPT focuses on the reactive segments in the trajectory of $X_{n}$. Specifically, a trajectory segment $X_{m},X_{m+1},\dots,X_{m+j}$, where $m\geq 0$ and $j\geq 1$, is called reactive, if it starts from a state in $A$ and it goes to $B$ without returning to $A$, or equivalently, if $X_{m}\in A$, $X_{m+j}\in B$, and $X_{i}\in(A\cup B)^{c}$ for all $m<i<m+j$. See Figure 2 for illustration. Suppose that an infinitely long trajectory of $X_{n}$ is given and let $M_{N}^{R}$ denote the number of reactive segments that occur within the first $N$ steps of the trajectory (the last reactive segment is allowed to end after the $N$th step). Then, the transition rate of the process $X_{n}$ from $A$ to $B$ is defined as
@@
Let us also recall the so-called (forward) committor $q$ associated to the sets $A$ and $B$, which is defined by
$q(x)=\mathbb{P}(\text{starting from $x$, $X_{n}$ enters $B$ before it enters $A$}),\quad x\in\mathbb{R}^{d}\,,$ (32)
and satisfies the equation
@@
    $k_{AB}=\int_{A}\Big{[}\int_{\mathbb{R}^{d}}p(x,y)q(y)\,\mathrm{d}{y}\Big{]}\pi(x)\,\mathrm{d}{x}=\int_{B}\Big{[}\int_{\mathbb{R}^{d}}p(x,y)(1-q(y))\,\mathrm{d}{y}\Big{]}\pi(x)\,\mathrm{d}{x}$.
@@
    $k_{AB}=\mathcal{E}(q)=\frac{1}{2}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}p(x,y)|q(y)-q(x)|^{2}\pi(x)\,\mathrm{d}{x}\,\mathrm{d}{y}$.
Proof.
$M_{N}^{R}=\sum_{m=0}^{N-1}\mathbbm{1}_{A}(X_{m})\mathbbm{1}_{B}(X_{m+j_{m}})\,,$
$k_{AB}=\lim_{N\rightarrow+\infty}\frac{1}{N}\sum_{m=0}^{N-1}\mathbbm{1}_{A}(X_{m})\mathbbm{1}_{B}(X_{m+j_{m}})=\int_{A}\Big{[}\int_{\mathbb{R}^{d}}p(x,y)q(y)\,\mathrm{d}{y}\Big{]}\pi(x)\,\mathrm{d}{x}\,,$
where the second equality follows from the definition of the committor $q$ in (32) and the ergodicity of the process $X_{n}$. Similarly, by considering reactive segments from $B$ to $A$ and applying the identity in the first claim, we obtain the second equality in this claim, since
$=$ $\int_{B}\Big{[}\int_{\mathbb{R}^{d}}p(x,y)(1-q(y))\mathrm{d}{y}\Big{]}\pi(x)\,\mathrm{d}{x}\,.$
Concerning the third claim, from (11) we can compute
$\mathcal{E}(q)=$ $\frac{1}{2}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}p(x,y)|q(y)-q(x)|^{2}\pi(x)\,\mathrm{d}{x}\,\mathrm{d}{y}$
$=$ $\langle q,(\mathcal{I}-\mathcal{T})q\rangle_{\mu}$
$=$ $\int_{A}q(x)[(\mathcal{I}-\mathcal{T})q](x)\pi(x)\mathrm{d}{x}+\int_{(A\cup B)^{c}}q(x)[(\mathcal{I}-\mathcal{T})q](x)\pi(x)\,\mathrm{d}{x}$
$+\int_{B}q(x)[(\mathcal{I}-\mathcal{T})q](x)\pi(x)\,\mathrm{d}{x}$
$=$ $\int_{B}[(\mathcal{I}-\mathcal{T})q](x)\pi(x)\,\mathrm{d}{x}$
$=$ $\int_{B}\pi(x)\,\mathrm{d}{x}-\int_{B}\Big{[}\int_{\mathbb{R}^{d}}p(x,y)q(y)\,\mathrm{d}{y}\Big{]}\pi(x)\,\mathrm{d}{x}$
$=$ $\int_{B}\Big{[}\int_{\mathbb{R}^{d}}p(x,y)(1-q(y))\,\mathrm{d}{y}\Big{]}\pi(x)\,\mathrm{d}{x}$
$=$ $k_{AB}\,,$
where the second equality follows from (12) in Lemma 13 and the fact that $\langle f,(\mathcal{I}-\mathcal{T}^{rev})g\rangle_{\mu}=\langle f,(\mathcal{I}-\mathcal{T})g\rangle_{\mu}$ when $f=g$, the fourth equality follows by using the equation (33) of the committor $q$, the fifth equality follows from the definition of $\mathcal{T}$ and $f|_{B}\equiv 1$, the sixth equality follows from the fact that $\int_{\mathbb{R}^{d}}p(x,y)\mathrm{d}{y}=1$, and the last equality follows from (34). ∎
It is worth mentioning that Proposition 3 also holds for nonreversible ergodic Markov processes. In particular, the third claim of Proposition 3 states that the transition rate $k_{AB}$ equals to the Dirichlet energy of the committor $q$. In the reversible setting (see (9)), we can further obtain the following variational characterization of the committor.
@@
Assume that $X_{n}$ is reversible. Let $q$ be the committor (32) associated to the transitions of $X_{n}$ from $A$ to $B$, and let $\mathcal{E}$ denote the Dirichlet energy defined in (11). For any function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ satisfying
$f|_{A}\equiv 0,\quad f|_{B}\equiv 1,$ (35)
@@
$\mathcal{E}(f)=k_{AB}+\mathcal{E}(f-q)\,.$ (36)
@@
$\min_{f}\mathcal{E}(f)$ (37)
among all functions $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ satisfying the condition (35), and the minimum value is $k_{AB}$.
@@
where the first, the third, and the fourth equalities follow from (12) in Lemma 13 and the self-adjointness of $\mathcal{T}$, the last equality follows from the third claim of Proposition 3, the fact that $f\equiv q$ on $A\cup B$ thanks to the condition (35) (which is met by both $f$ and $q$), as well as the identity $(\mathcal{I}-\mathcal{T})q=0$ on $(A\cup B)^{c}$ as stated in the equation (33). ∎
Similar to Theorem 2.3 which concerns the leading eigenfunctions, we have the following characterization of the committor as a slow variable for the study of transitions from $A$ to $B$. We also refer to Remark 4.2 for discussions on a numerical approach (based on Proposition 4 and the effective dynamics in the next section) for learning CV maps that can parametrize the committor.
@@
$\min_{f}\lim_{N\rightarrow+\infty}\frac{1}{2N}\sum_{n=0}^{N-1}|f(X_{n+1})-f(X_{n})|^{2}$
among all functions $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ which satisfy $f|_{A}\equiv 0$ and $f|_{B}\equiv 1$.

### v2

@@ 2.2 Spectrum and timescales
In this section, we study the spectrum of $\mathcal{T}$ as well as its implications on the speed of convergence of the dynamics $X_{n}$ to equilibrium.
We assume that $X_{n}$ is reversible. By (9), $\mathcal{T}$ is self-adjoint with respect to the inner product (3), and therefore its spectrum $\sigma(\mathcal{T})$ is real. To simplify the discussion, we further confine ourself to the case where each nonzero element of $\sigma(\mathcal{T})$ is an eigenvalue $\lambda\in\mathbb{R}$, for which the (integral) equation
@@
The proof of Proposition 1 is presented in Appendix B. The next proposition gives conditions which guarantee that all eigenvalues are nonnegative.
@@
    There is a unique self-adjoint linear operator $\mathcal{A}$ on $\mathcal{H}$ such that $\mathcal{A}\geq 0$ and $\mathcal{T}=\mathcal{A}^{2}$.
@@
Note that by the definition of $\mathcal{T}$ in (2) we have
$\begin{split}\langle\mathcal{T}f,f\rangle_{\mu}=&\int_{\mathbb{R}^{d}}\Big{(}\mathbb{E}(f(X_{1})|X_{0}=x)\Big{)}f(x)\mu(\mathrm{d}{x})=\mathbb{E}_{X_{0}\sim\mu}(f(X_{0})f(X_{1}))\,,~{}f\in\mathcal{H}\,,\end{split}$ (22)
where the second equality follows from the law of total expectation. Therefore, the equivalence between (a) and (b) follows from the variational principle for the smallest eigenvalue of $\mathcal{T}$ (see [56, Theorem 2.20]). To prove the equivalence between (b) and (c), we notice that (22) implies that (b) is equivalent to $\mathcal{T}\geq 0$. Also, it is easy to verify that $\mathcal{T}$ is bounded. Therefore, applying the square root lemma [51, Theorem VI.9] we see that (b) implies (c). Conversely, assuming (c), we have $\langle\mathcal{T}f,f\rangle_{\mu}=\langle\mathcal{A}^{2}f,f\rangle_{\mu}=\langle\mathcal{A}f,\mathcal{A}f\rangle_{\mu}\geq 0$ and therefore (b) holds. This shows that (b) and (c) are equivalent. ∎
@@
Roughly speaking, the condition (b) of Proposition 2 describes that “on average, $f(X_{n})$ does not change sign in two consecutive steps”.
A situation under which the condition (c) is met is when the chain $X_{n}$ can be embedded into another reversible Markov process $Y_{m}$, such that $X_{n}=Y_{2n}$ for all $n\geq 0$. In particular, suppose that $X_{n}$ corresponds to the state of a reversible diffusion process governed by a SDE with a generator $\mathcal{L}$ and let $\tau>0$ be the lag-time (see the first case in Example 2.1). Then, $\mathcal{T}=\mathrm{e}^{\tau\mathcal{L}}=\mathcal{A}^{2}$ with $\mathcal{A}=\mathrm{e}^{\frac{\tau}{2}\mathcal{L}}$. In fact, in this case all eigenvalues of $\mathcal{T}$ are positive.
In general, it is possible that a transfer operator $\mathcal{T}$ has negative eigenvalues $\lambda\in(-1,0)$ under the assumptions in Proposition 1. Such negative eigenvalues (and the corresponding eigenfunctions) may even be important in studying the large-time dynamics of $X_{n}$, if they are close to $-1$. Similar to Theorem 2.2 below, variational characterization for those negative eigenvalues (which are close to $-1$) can be obtained and applied to developing numerical methods. However, we will not consider negative eigenvalues, since we are mainly interested in the case where the Markov chains come from reversible SDEs and the eigenvalues in this case are all positive according to the discussion above.
@@
Moreover, by Hilbert-Schmidt theorem [51, Theorem VI.16], we can choose the corresponding eigenfunctions $(\varphi_{i})_{i\geq 0}$, with $\varphi_{0}\equiv 1$, such that they form an orthonormal basis of $\mathcal{H}$. For a function $f\in\mathcal{H}$, from (2) we obtain
@@
where $\mathbb{E}_{\mu}(f)$ denotes the mean value of $f$ under $\mu$. Therefore, as the Markov chain evolves, the expectation $\mathbb{E}(f(X_{n})|X_{0}=x)$ converges to $\mathbb{E}_{\mu}(f)$ (i.e. the value at equilibrium), and the speed of convergence is determined by the leading eigenvalues of $\mathcal{T}$ that are closest to $1$. Estimating these leading eigenvalues and computing the corresponding eigenfunctions help to understand the long-time dynamical behaviors (e.g. identifying large timescales and metastable states) of the process $X_{n}$. In this regard, let us present the following variational characterization of the large eigenvalues and the corresponding eigenfunctions of $\mathcal{T}$ (see [69, Theorem 2.1]).
@@
$\sum_{i=1}^{m}\omega_{i}(1-\lambda_{i})=\min_{f_{1},\dots,f_{m}\in\mathcal{H}}\sum_{i=1}^{m}\omega_{i}\mathcal{E}(f_{i})\,,$ (25)
@@
$\mathbb{E}_{\mu}(f_{i})=0,\quad\langle f_{i},f_{j}\rangle_{\mu}=\delta_{ij}\,,\quad\forall\,i,j\,\in\{1,\dots,m\}\,.$ (26)
@@
We mention that similar variational characterizations for eigenvalues of generator have been obtained in [68, Theorem 3] and [67, Theorem 1]. The proof of Theorem 2.2, which we present in Appendix B, is adapted from the proof of [67, Theorem 1, in Appendix A].
In the following remark we discuss the connection between Theorem 2.2 and the variational formulations in VAMP [61]. Despite the similarities, we emphasize that, by employing the Dirichlet energy $\mathcal{E}$, we are able to draw a connection between eigenfunctions and slow CVs using Theorem 2.2 (see Theorem 2.3 and the discussions after Remark 2.3), and obtain a unified framework that characterizes both the spectrum and transition rates (see Proposition 4 in Section 2.3) of the process.
@@
Theorem 2.2 is similar to the variational formulation in VAMP in the reversible case (see [61, Appendix D] and [63, Section II.A]). In fact, using the identity (12) in Lemma 13 and the normalization constraints in (26), we can write (25) as
$\sum_{i=1}^{m}\omega_{i}\lambda_{i}=\max_{f_{1},\dots,f_{m}\in\mathcal{H}}\sum_{i=1}^{m}\omega_{i}\langle f_{i},\mathcal{T}f_{i}\rangle_{\mu}\,,$ (27)
which recovers the VAMP-1 score defined in [61, Theorem 2] when $\omega_{i}=1$ for $1\leq i\leq m$ (the eigenvalue $\lambda_{0}=1$ and its eigenfunction are excluded by imposing the first constraint in (26)).
@@
To explain the idea, let us first assume $k=1$ and consider scalar CV maps. Our observation is that, for a CV map $\xi:\mathbb{R}^{d}\rightarrow\mathbb{R}$ to be indicative of essential transitions of $X_{n}$ on large timescales, it is necessary that $\xi$ is insensitive to the noisy fluctuations on the smallest timescales. In other words, $\xi$ should be a slow variable such that the observations $\xi(X_{0}),\xi(X_{1}),\dots$ vary slowly (in most of the time unless essential transitions occur) as the dynamics $X_{n}$ evolves. This reasoning suggests that a natural way to find a good CV map $\xi$ is by minimizing the average of squared variations
@@
The above reasoning can be extended to the case of vector-valued CV maps. Let us summarize it in the following result.
@@
Assume that $\omega=(\omega_{1},\dots,\omega_{m})^{\top}\in\mathbb{R}^{m}$ has positive components, where $\omega_{1}\geq\omega_{2}\geq\dots\geq\omega_{m}>0$. Denote by $|z|_{\omega}:=(\sum_{i=1}^{m}\omega_{i}|z_{i}|^{2})^{1/2}$ the weighted $l^{2}$-norm of vectors $z=(z_{1},\dots,z_{m})^{\top}\in\mathbb{R}^{m}$. The CV map $\xi=(\varphi_{1},\dots,\varphi_{m})^{\top}$ defined by the first eigenfunctions $\varphi_{1},\dots,\varphi_{m}$ minimizes the average of squared variations
@@
among functions $f=(f_{1},\dots,f_{m})^{\top}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ whose components belong to $\mathcal{H}$ and satisfy (26).
@@
Using the identity (13) in Lemma 13, we can derive the objective in (29) as
$\lim_{N\rightarrow+\infty}\frac{1}{2N}\sum_{n=0}^{N-1}|f(X_{n+1})-f(X_{n})|^{2}_{\omega}$
$=$ $\sum_{i=1}^{m}\omega_{i}\left(\lim_{N\rightarrow+\infty}\frac{1}{2N}\sum_{n=0}^{N-1}|f_{i}(X_{n+1})-f_{i}(X_{n})|^{2}\right)=\sum_{i=1}^{m}\omega_{i}\mathcal{E}(f_{i})\,.$
Therefore, Theorem 2.2 implies that the minimum of (29) under the constraints (26) is achieved when $f$ is defined by the first $m$ eigenfunctions. ∎
Note that the norm in (29) is an extension of the norm in (28) when $k>1$. The weights $\omega$ are introduced in (29) in order to remove the non-uniqueness of the minimizer of (29) due to reorderings of $f_{1},f_{2},\dots,f_{m}$. Condition (26) guarantees that $f_{1},f_{2},\dots,f_{m}$ are non-constant, normalized, and pairwise orthogonal (therefore linearly independent).
@@ 2.3 Transition rates
In this section, we study the transition of the (discrete-in-time and continuous-in-space) process $X_{n}$ between two subsets $A,B\subset\mathbb{R}^{d}$ in the framework of TPT [57, 15, 16]. We refer to [59, 21, 20] for extensions of TPT to different types of Markov chains.
Let us assume that $A,B\subset\mathbb{R}^{d}$ are two closed subsets with smooth boundaries $\partial A$ and $\partial B$, respectively, such that $A\cap B=\emptyset$ and the complement of their union is nonempty, i.e. $(A\cup B)^{c}:=\mathbb{R}^{d}\setminus(A\cup B)\neq\emptyset$. Recall that TPT focuses on the reactive segments in the trajectory of $X_{n}$. Specifically, a trajectory segment $X_{m},X_{m+1},\dots,X_{m+j}$, where $m\geq 0$ and $j\geq 1$, is called reactive, if it starts from a state in $A$ and it goes to $B$ without returning to $A$, or equivalently, if $X_{m}\in A$, $X_{m+j}\in B$, and $X_{i}\in(A\cup B)^{c}$ for all $m<i<m+j$. See Figure 2 for illustration. Denote by $M_{N}^{R}$ the number of reactive segments that occur within the first $N$ steps of the trajectory (the last reactive segment is allowed to end after the $N$th step). Then, the transition rate of the process $X_{n}$ from $A$ to $B$ is defined as
@@
Let us also recall the (forward) committor $q$ associated to the sets $A$ and $B$, which is defined by
$q(x)=\mathbb{P}(\text{starting from $x$, $X_{n}$ enters $B$ before it enters $A$}),\quad x\in\mathbb{R}^{d}\,.$ (31)
In mathematical literatures, the committor $q$ is called equilibrium potential, which plays a central role in the study of metastability [9, 10, 5]. Calculating the probability in (31) by considering the state in the next step, it is straightforward to verify that $q$ satisfies the equation (see [42, Appendix A] for the proof in the case of Markov jump processes)
@@
    $k_{AB}=\int_{A}\left[\int_{\mathbb{R}^{d}}p(x,y)q(y)\,\mathrm{d}{y}\right]\mu(\mathrm{d}{x})=\int_{B}\left[\int_{\mathbb{R}^{d}}p(x,y)(1-q(y))\,\mathrm{d}{y}\right]\mu(\mathrm{d}{x})$.
@@
    $k_{AB}=\mathcal{E}(q)=\frac{1}{2}\int_{\mathbb{R}^{d}}\left[\int_{\mathbb{R}^{d}}p(x,y)(q(y)-q(x))^{2}\mathrm{d}{y}\right]\mu(\mathrm{d}{x})$.
The proof of Proposition 3 is presented in Appendix B. It is worth mentioning that the proof of Proposition 3 does not rely on reversibility and hence Proposition 3 holds for non-reversible ergodic Markov processes as well. In particular, the third claim of Proposition 3 relates the transition rate $k_{AB}$ to the Dirichlet energy of the committor $q$. In the reversible setting (see (9)), we can further obtain the following variational characterization of the committor.
@@
Assume that $X_{n}$ is reversible. Let $q$ be the committor (31) associated to the transitions of $X_{n}$ from $A$ to $B$, and let $\mathcal{E}$ denote the Dirichlet energy defined in (11). Define the space
$\mathcal{F}_{AB}=\Big{\{}f\in\mathcal{H}\,\Big{|}\,f|_{A}\equiv 0,~{}f|_{B}\equiv 1\Big{\}}\,.$ (33)
@@
$\mathcal{E}(f)=k_{AB}+\mathcal{E}(f-q)\,,\quad\forall~{}f\in\mathcal{F}_{AB}\,.$ (34)
@@
$\min_{f\in\mathcal{F}_{AB}}\mathcal{E}(f)\,,$ (35)
and the minimum value of (35) is $k_{AB}$.
@@
where the first, the third, and the fourth equalities follow from (12) in Lemma 13 and the self-adjointness of $\mathcal{T}$, the last equality follows from the third claim of Proposition 3, the fact that $f\equiv q$ on $A\cup B$ thanks to the boundary condition in (33) (which is met by both $f$ and $q$), as well as the identity $(\mathcal{I}-\mathcal{T})q=0$ on $(A\cup B)^{c}$ as stated in the equation (32). ∎
We note that the variational formulation (35) is well-known in the literatures on TPT [16] and on metastability [9].
Similar to Theorem 2.3 which concerns the leading eigenfunctions, we have the following characterization of the committor as a slow variable for the study of transitions from $A$ to $B$.
@@
$\min_{f\in\mathcal{F}_{AB}}\lim_{N\rightarrow+\infty}\frac{1}{2N}\sum_{n=0}^{N-1}|f(X_{n+1})-f(X_{n})|^{2}\,,$
where $\mathcal{F}_{AB}$ is the space defined in (33).